Main INDEX
Monthly INDEX
PREV
NEXT
Make New Entry,
Make Followup Entry
User name riordan
Log entry time 21:47:06 on March 02, 2009
Entry number 264237
keyword=Summary for March 1 Swing - March 2 Day
Summary for Mar 1 Swing - Mar 2 Day
Practically no data taking due to CHL recovery and cryomodule thermal
cycling
Beam in Use
Shift |
ABU |
BANU |
BNA |
2009/03/01 Swing | 3.05 |
0.30 | 4.65 |
2009/03/02 Owl | 0.00 |
0.00 | 8.00 |
2009/03/02 Day | 0.00 |
0.00 | 8.00 |
Total | 3.05 |
0.30 | 20.65 |
Main events/issues:
- March 1
- Beam quality continues to degrade. Very little usable data is
taken.
- Around 23:00, power glitches trip the CHL
- These glitches also cause problems in the HRS, in particular
we lose control to an IOC that controls the HV
- March 2
- No beam while MCC is recovering
- Replaced UPS powering network switch for HRS HV IOC.
- Left HRS put in positive polarity for positron runs
- No beam anticipated until Owl 3/3
Plans for the next couple of days:
- March 3 - Run through 5-pass positron kinematics
- Moller measurment after HRS polarity switch back to electron
- Take as much 5-pass production with goals of 50% complete data
points until pass change
- Pass change will depend on the outcome of an appeal for extended
running to be determined tomorrow afternoon. Without extension, we plan
on changing Wednesday morning
A copy of this log entry has been emailed to: brads, jpchen, jiang, meziani, choi